# DimOL: Dimensional Awareness as A New ‘Dimension’ in Operator Learning

### v1

@@ 3.4 Transformer-Based Models with ProdLayer
The Latent Spectral Method (LSM) is a cutting-edge Transformer-based model designed for solving PDEs. Its basic innovation is the incorporation of Sine functions into the network, allowing it to approximate any (periodic) function within a specific spectral range. For non-periodic components, LSM utilizes convolution-based neural blocks. The overall structure of LSM is akin to a U-Net, enabling it to learn interactions between different patches of the domain at varying scales.
Inspired by the separation of variables method, we hypothesize that the projection in the latent spectral space can be enhanced if terms can multiply Sine or Cosine functions directly, rather than merely adding them as biases. This modification aligns with the core principle of DimOL, which aims to improve model performance by encoding product terms. To implement this idea, we replace the linear transformation in the decoder of LSM with a ProdLayer, as shown in Figure1(c). This enhanced architecture retains the original strengths of LSM while integrating the ProdLayer to facilitate improved spectral projections, allowing the model to handle product terms more effectively. Given this intuition, a larger number of product terms for each ProdLayer is beneficial and can enhance the model’s capacity to capture complex interactions.
@@ 4.1 Experimental Setups
Hyperparameter.
The following experiments use a product number of $p=2$ in each ProdLayer (see Figure 1). While this empirical value may not be the optimal solution for every specific dataset, it consistently leads to improvements in our experiments. Other options for the $p$ value also yield performance improvements and are discussed in the ablation study in Section 4.8. Specifically, setting $p=0$ would reduce the channel-mixing layer of the model to an ordinary MLP. For FNO and LSM models, we adopt the original depth settings of $4$ layers.
Table 1: Results on the Burgers dataset, where chan. mix denotes the use of a channel mixing layer.
Model # Parameters (Million) MSE (×$\times 10^{-3}$) Gain
FNO w/o chan. mix (Li et al., 2021) 0.098 0.646 -
FNO w/ chan. mix 0.106 1.045 -61.8%
FNO + ProdLayer 0.105 0.593 8.2%
T-FNO w/o chan. mix (Li et al., 2021) 0.052 2.225 -
T-FNO w/ chan. mix 0.056 2.342 -5,3%
T-FNO + ProdLayer 0.056 1.147 48.4%
@@ 4.2 Burgers
We adopt the Burgers dataset of the FNO paper. The 1D Burgers’ equation is given by $\partial_{t}u(x,t)+u(x,t)\partial_{x}(u(x,t))=\nu\partial_{xx}u(x,t),\ x\in(0,1).t\in(0,1]$. Since this is a 1D problem, more complex methods like F-FNO, T-FNO, and LSM are not applicable, so we compare FNO with its dimension-aware counterpart, FNO + ProdLayer. As shown in Table 1, we evaluate the $4$-layer T-FNO models on the Burgers dataset and achieve an impressive $48.4\%$ drop in MSE on the test set. We also compare the model with the version with channel-mixing of MLP, which would overfit and have lower performance. In contrast, the models with ProdLayers as channel-mixing are less likely to overfit, due to its intrinsic dimensional awareness.
@@ 4.3 Darcy Flow
This dataset is also sourced from FNO. The physical process of Darcy Flow can be described by the equation $-\nabla\cdot(a(x)\nabla u(x))=f(x),\ \ x\in(0,1)^{2}$, where the input is $a(x)$ and the output is $u(x)$, which satisfies the steady-state equation. It’s important to note that this setting focuses on finding a steady state under given conditions, differing from dynamic prediction tasks. As shown in Table 2, the DimOL-adapted models consistently outperform the original versions.
Table 2: Model performance on the Darcy Flow dataset. Notably, the use of DimOL provides a consistent performance improvement to all compared models.
Model # Parameters (Million) MSE (×$\times 10^{-3}$) Gain
FNO (Li et al., 2021) 1.658 5.889 -
FNO + ProdLayer 1.663 5.726 2.8%
T-FNO (Kossaifi et al., 2023) 0.684 7.500 -
T-FNO + ProdLayer 0.688 6.919 7.8%
LSM (Wu et al., 2023) 4.813 2.892 -
LSM + ProdLayer 4.812 2.713 6.2%
@@ 4.4 Navier-Stokes
Results on TorusLi.
As shown in Table 3, DimOL models consistently outperform the base models on the TorusLi dataset. Additionally, we test the product term encoding with quadratic layers (i.e., ProdLayer). However, as demonstrated in the tests on FNO and T-FNO, the quadratic layer does not surpass the performance of the more straightforward ProdLayer.
Table 3: Model performance on the TorusLi dataset.
      Base model       DimOL       MSE       Gain
      FNO (Li et al., 2021)       -       0.1496       -
      FNO       QuadLayer       0.1431       4.3%
      FNO       ProdLayer       0.1394       6.8%
      T-FNO (Kossaifi et al., 2023)       -       0.1448       -
      T-FNO       QuadLayer       0.1430       1.2%
      T-FNO       ProdLayer       0.1404       2.3%
      F-FNO (Tran et al., 2023)       -       0.1651       -
      F-FNO       ProdLayer       0.1542       6.6%
Results on TorusVisForce.
Due to the size of the TorusVisForce dataset, we select T-FNO for our case study, given its strong performance on TorusLi. In the TorusVisForce dataset, the involvement of more physical quantities in the input allows the DimOL models to achieve significant improvements. We conduct two tasks: in the first, we set the prediction horizon to $T=4$ time steps, while in the second, we have $T=10$. Obviously, as the prediction horizon extends, the task becomes more challenging. Table 4 presents the prediction errors. We can observe that our DimOL method significantly improves the performance of the base models in both tasks.
Table 4: Model performance on the TorusVisForce dataset with different prediction horizons.
Model $T=4$ $T=10$
MSE (×$\times 10^{-2}$) Gain MSE (×$\times 10^{-2}$) Gain
T-FNO (Kossaifi et al., 2023) 1.090 - 1.717 -
T-FNO + ProdLayer 0.966 11.5% 1.622 5.5%
@@ 4.5 Inductor2D
Table 5: Prediction errors on the Inductor2D dataset with irregular meshes.
Model MSE-$A_{z}$ (×$\times 10^{-2}$) MSE-$B_{x}$ MSE-$B_{y}$ Gain
GNOT (Hao et al., 2023) 0.8736 0.2579 0.3873 -
GNOT + ProdLayer 0.8869 0.2578 0.3847 4.8%
4.6 Few-Shot Learning Studies on Operator Properties
We aim to develop a model that approximates an operator while preserving key properties of the original operator. If successful, the model should generalize well in few-shot learning scenarios. To evaluate this, we design three experimental settings:
  • Datasets: We construct two subsets of the TorusVisForce dataset, named TorusVisForce-TopMu and -LowMu. These datasets contain $200$ and $100$ data sequences respectively with the highest and lowest viscosity values ranging in $[8.2\times 10^{-5},1\times 10^{-4}]$ and $[1\times 10^{-5},1.8\times 10^{-5}]$.
  • Few-shot training: We use $100$ random samples from TorusVisForce-TopMu as the training set.
  • In-set testing: We first perform in-distribution testing on the other partition of $100$ unseen training samples from the TorusVisForce-TopMu subset.
  • Partial-OOD testing: We generate a partially OOD test set by scaling the training data using a coefficient $k$. Here, $k$ is an integer applied to the physical quantities $\omega$, $f$, and $\nu$.
  • OOD testing: We finally use TorusVisForce-LowMu as the test set to evaluate the model’s generalization ability to OOD data.
For Partial-OOD testing, we here leverage a crucial principle from dimensional analysis: the invariance of similar transformations. This principle motivates us to create partially out-of-distribution test data using simple data transformations to assess operator learning. The transformed data, while indeed outside the original data distribution, effectively maintains similar dynamics to the training data. We use the TorusVisForce dataset as a case study. Consider data transformations:
@@ 4.6 Few-Shot Learning Studies on Operator Properties
which is validated using a traditional numerical solver, although a slight numerical bias is observed. Therefore, in the Partial-OOD experiments, although DimOL cannot access the analytic expression for the actual operator $G^{\dagger}:[H_{per}^{r}((0,1)^{2}),L_{per}^{2}((0,1)^{2};\mathbb{R}),\mathbb{R}_{+}]\rightarrow H_{per}^{r}((0,1)^{2})$ in TorusVisForce, it will seek to uncover the key properties from the transformed data.
Table 6: Prediction errors (×$\times 10^{-2}$) on TorusVisForce ($T=4$) in few-shot learning and OOD setups.
Base Method MSE (In-Set) MSE ($k=4$) MSE ($k=16$ ) MSE (OOD)
T-FNO w/o chan. mix 2.354 1.158 1.246 8.975
T-FNO $p=0$ 1.625 1.005 1.062 8.000
T-FNO $p=2$ 1.456 0.911 0.984 7.828
Gain 10.4% 9.4% 7.3% 2.2%
LSM $p=0$ 8.304 1.427 1.500 16.390
LSM $p=2$ 8.031 1.262 1.334 16.040
Gain 3.3% 11.6% 11.1% 2.1%
@@
4.7 The Potential of Finding New Physical Terms with DimOL
Figure 2: The spectral terms regressed.
@@ 4.7 The Potential of Finding New Physical Terms with DimOL
$\frac{d}{dx}g(x)=\mathcal{F}^{-1}[ik(\mathcal{F}g(x))];\quad\frac{d^{2}}{dx}g(x)=\mathcal{F}^{-1}[-k^{2}(\mathcal{F}g(x))].$ (9)
@@
4.8 Hyperparameter Sensitivity Analysis
@@
Table 7: Sensitivity analysis on various datasets for the hyperparameter choice of $p$ used in ProdLayer. Results are presented in MSE (×$\times 10^{-3}$). For TorusLi, we set the prediction horizon to $T=10$. For TorusVisForce, we conduct the few-shot experiments with $N=100$ and $T=4$. Results marked with an underline indicate the worst performance, while those in bold represent the best.
Dataset Base w/o chan. mix $p=0$ $p=1$ $p=2$ $p=4$ $p=8$ $p=16$
@@
5 Conclusions and Limitations
In this paper, we introduced DimOL as an adaptation for major models tailored to solving high-dimensional PDEs. These include FNO-based models and Transformer-based models such as LSM and GNOT. We discovered that dimensional awareness not only makes sense for human physicists but also machine learning models, especially when the task involves actual physical quantities as input. Our research indicates that this conclusion applies across a wide range of models and datasets. By enhancing the model’s symbolic meaning, DimOL models are better suited for reverse engineering to uncover terms with potential physical significance. While we currently use ProdLayer for DimOL, we believe that DimOL is more than this. It should be regarded as a new “dimension” to consider when designing AI-for-Science models.
A potential limitation of DimOL could be its effectiveness in tasks where the input data lacks clear dimensional information or where the relationship between dimensions is highly complex or nonlinear. In such cases, DimOL may not offer significant improvements.
Acknowledgememts
This work was supported by the National Natural Science Foundation of China (Grant No. 62250062, 62106144), the Shanghai Municipal Science and Technology Major Project (Grant No. 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.

### v2

@@ 3.4 Transformer-Based Models with ProdLayer
LSM is a cutting-edge Transformer-based model designed for solving PDEs. Its basic innovation is the incorporation of Sine functions into the network, allowing it to approximate any (periodic) function within a specific spectral range. For non-periodic components, LSM utilizes convolution-based neural blocks. The overall structure of LSM is akin to a U-Net, enabling it to learn interactions between different patches of the domain at varying scales.
Inspired by the separation of variables method, we hypothesize that the projection in the latent spectral space can be enhanced if terms can multiply Sine or Cosine functions directly, rather than merely adding them as biases. This modification aligns with the core principle of DimOL, which aims to improve model performance by encoding product terms. To implement this idea, we replace the linear transformation in the decoder of LSM with a ProdLayer, as shown in Figure1(c). This enhanced architecture retains the original strengths of LSM while integrating ProdLayer allows the model to handle product terms more effectively. Given this intuition, a larger number of product terms for each ProdLayer is beneficial and can enhance the model’s capacity to capture complex interactions.
@@ 4 Experiments
Table 1: Prediction results on the Burgers dataset. We use ProdLayer in the DimOL methods and show that our approach consistently improves the performance of the base models.
Model # Param MSE ($10^{-3}$) Gain
FNO (2021) 0.098M 1.683 -
FNO + DimOL 0.105M 1.428 9.3%
T-FNO (2021) 0.052M 2.225 -
T-FNO + DimOL 0.056M 1.147 48.4%
CNO (2024) 1.928M 73.255 -
CNO + DimOL 1.912M 64.684 11.7%
@@ 4.1 Experimental Setups
The following experiments use a product number of $p=2$ in each ProdLayer (see Figure 1). While this empirical value may not be the optimal solution for every specific dataset, it consistently leads to improvements in our experiments. Other options for the $p$ value also yield performance improvements and are discussed in Appendix B. Specifically, setting $p=0$ would reduce the channel-mixing layer of the model to an ordinary MLP. For FNO and LSM models, we adopt the original depth settings of $4$ layers.
@@ 4.2 Burgers
We adopt the 1D Burgers dataset of the FNO paper, which can be described using the following equation: $\partial_{t}u(x,t)+u(x,t)\partial_{x}(u(x,t))=\nu\partial_{xx}u(x,t),\ x\in(0,1).t\in(0,1]$. Since this is a 1D problem, more complex methods like F-FNO, T-FNO, and LSM are not applicable, so we compare FNO with its dimension-aware counterpart, FNO + ProdLayer. As shown in Table 1, we evaluate the $4$-layer T-FNO models on the Burgers dataset and achieve an impressive $48.4\%$ drop in MSE on the test set, which performs the best; the adapted FNO model also achieves a gain of $9.3\%$ on Burgers dataset. Similar effects also hold for CNO models.
Table 2: Prediction results on the Darcy Flow dataset. The use of DimOL provides consistent performance gains.
Model # Param MSE ($10^{-3}$) Gain
FNO (2021) 1.658M 5.889 -
FNO + DimOL 1.663M 5.726 2.8%
T-FNO (2023) 0.684M 7.500 -
T-FNO + DimOL 0.682M 6.919 7.8%
LSM (2023) 4.813M 2.892 -
LSM + DimOL 4.812M 2.713 6.2%
CNO  (2024) 5.582M 9.525 -
CNO + DimOL 5.578M 8.773 7.9%
Table 3: Prediction results on the TorusLi dataset. We conduct experiments with alternative architectures to implement DimOL.
Base model DimOL MSE Gain
FNO - 0.1496 -
FNO QuadLayer 0.1431 4.3%
FNO ProdLayer 0.1394 6.8%
T-FNO - 0.1448 -
T-FNO QuadLayer 0.1430 1.2%
T-FNO ProdLayer 0.1404 2.3%
F-FNO - 0.1651 -
F-FNO ProdLayer 0.1542 6.6%
CNO - 0.2044 -
CNO ProdLayer 0.1893 7.4%
Table 4: Prediction results (×$\bf\times 10^{-2}$) on the TorusVisForce dataset. $T$ denotes different prediction time horizons.
Model $T=4$ $T=10$
MSE Gain MSE Gain
T-FNO (2023) 1.090 - 1.717 -
T-FNO + DimOL 0.966 11.5% 1.622 5.5%
CNO (2024) 3.457 - 6.264 -
CNO + DimOL 3.308 4.3% 5.838 6.8%
@@ 4.3 Darcy Flow
This dataset is also sourced from FNO. The physical process of Darcy Flow can be described by the equation $-\nabla\cdot(a(x)\nabla u(x))=f(x),\ \ x\in(0,1)^{2}$, where the input is $a(x)$ and the output is $u(x)$, which satisfies the steady-state equation. Notably, this setting focuses on finding a steady state under given conditions, differing from dynamic prediction tasks. As shown in Table 2, the DimOL-adapted models consistently outperform the original versions.
@@ 4.4 Navier-Stokes
TorusLi.
As shown in Table 3, DimOL models consistently outperform the base models on the TorusLi dataset. Additionally, we test the product term encoding with quadratic layers. However, as demonstrated in the tests on FNO and T-FNO, the quadratic layer does not surpass the performance of the more straightforward ProdLayer.
TorusVisForce.
Due to the size of the TorusVisForce dataset, we select T-FNO for our case study, given its strong performance on TorusLi. In the TorusVisForce dataset, the involvement of more physical quantities in the input allows the DimOL models to achieve significant improvements. We conduct two tasks: in the first, we set the prediction horizon to $T=4$ time steps, while in the second, we have $T=10$. Obviously, as the prediction horizon extends, the task becomes more challenging. Table 4 presents the prediction errors. We observe that DimOL significantly improves the performance of the base models in both tasks.
Table 5: MSE on the Inductor2D dataset with irregular meshes. We use GNOT as the baseline model here, as other compared methods did not include this dataset in their original papers.
Model $A_{z}$ ($10^{-2}$) $B_{x}$ $B_{y}$ Gain
GNOT (2023) 0.8736 0.2579 0.3873 -
GNOT + DimOL 0.8869 0.2578 0.3847 4.8%
@@ 4.5 Inductor2D
4.6 Out-of-Distribution Few-Shot Learning Studies
Table 6: OOD few-shot learning results in MSE (×$\bf\times 10^{-2}$) on TorusVisForce ($\bf T=4$). Here, $p$ is the hyperparameter in ProdLayer and $k$ denotes the scaling coefficient applied to the in-set physical quantities to construct OOD data.
Base DimOL In-Set OOD ($k=4$) OOD ($k=16$ )
T-FNO - 2.354 1.158 1.246
T-FNO $p=0$ 1.625 1.005 1.062
T-FNO $p=2$ 1.456 0.911 0.984
Gain 10.4% 9.4% 7.3%
LSM $p=0$ 8.304 1.427 1.500
LSM $p=2$ 8.031 1.262 1.334
Gain 3.3% 11.6% 11.1%
We aim to develop a model that approximates an operator while preserving the key properties of the original operator. If successful, the model should generalize well in few-shot learning scenarios. To evaluate this, we design the following experiments. We construct a subset of the TorusVisForce dataset, containing $200$ data sequences with viscosity ranging in $[8.2\times 10^{-5},1\times 10^{-4}]$. For few-shot learning, we use $100$ random samples from this dataset as the training set. We employ two test settings:
  • In-set test: We first perform in-distribution testing on the other partition of $100$ unseen samples.
  • OOD test: We generate an OOD test set by scaling the in-set test data using a coefficient $k$, where $k$ is an integer applied to the physical quantities $\omega$, $f$, and $\nu$.
For OOD testing, we leverage a crucial principle from dimensional analysis: the invariance of similar transformations. This principle motivates us to create partially out-of-distribution test data using simple data transformations to assess operator learning. The transformed data, while indeed outside the original data distribution, effectively maintains similar dynamics to the training data. We use the TorusVisForce dataset as a case study. Consider data transformations:
@@ 4.6 Out-of-Distribution Few-Shot Learning Studies
which is validated using a traditional numerical solver, although a slight numerical bias is observed. Therefore, in the OOD experiments, although DimOL cannot access the analytic expression for the actual operator $G^{\dagger}:[H_{per}^{r}((0,1)^{2}),L_{per}^{2}((0,1)^{2};\mathbb{R}),\mathbb{R}_{+}]\rightarrow H_{per}^{r}((0,1)^{2})$ in TorusVisForce, it will seek to uncover the key properties from the transformed data.
@@
4.7 Generalization to New Boundary Conditions
Another profound issue of machine-learning-based PDE solvers is how well they can generalize across different boundaries. In this work, we employ the cylinder-flow setting. In this setting, we consider cylinders of various shapes randomly placed within the domain, and we impose the Dirichlet Boundary Condition on the velocity field. The domain has open boundaries on the left and right sides, while the upper and lower sides act as closed boundaries. The initial condition specifies a zero-velocity field throughout the domain, except on the left side where the velocity is set to $v_{0}$. Soon after the fluid flow encounters the cylinder, the velocity field develops a pattern of Karman Vortex Street.
Our datasets consist of cylinders with different shapes: Shape C represents a circle, S represents a square, and T represents a triangle. We train the models on a training set with four types of boundary conditions: (1-3) 1C, 1S, and 1T, simulating flows with a single circle, square, and triangle, respectively; and (4) ST, which includes one square and one triangle. The test data includes cylinders of in-set boundaries 1C, 1S, 1T, and ST, as well as out-of-distribution (OOD) combinations of CS and CT. Given that reflections can occur between two cylinders, making accurate predictions on test sets with two-cylinder problems is challenging. However, since the model is trained on individual shapes C, S, and T, along with the combination ST, it is expected to generalize well to the unseen combinations CS and CT.
For this experiment, only FNO-based models are suitable for the rectangular domain. Therefore, we select FNO and T-FNO as our baseline models. As shown in Table 7, the DimOL models not only outperform the baseline models on the in-set datasets but also demonstrate excellent generalizability under unseen boundary conditions. A representative case is presented in Figure 2.
Figure 2: Heatmap of predicted velocity and corresponding errors under different boundary conditions, which can be different from those in the training data. We have $3$ types of cylinders: Circle, Square, and Triangle. (a): Task 1C — single Circle; (b): Task ST — one Square and one Triangle; (c): Task CT — one Circle and one Triangle. Notably, we train the models on four tasks: 1C, 1S, 1T, and ST. Therefore, (a) and (b) represent in-set test cases, while (c) represents OOD samples.
4.8 Finding New Physical Terms with DimOL
Figure 3: Results of regressed spectral terms.  (a) The imaginary part of the 1st Fourier weight, i.e., the linear term.  (b) The real part of the 2nd Fourier weight, i.e., the quadratic term.
@@ 4.8 Finding New Physical Terms with DimOL
$\frac{d}{dx}g(x)$ $=\mathcal{F}^{-1}[ik(\mathcal{F}g(x))],$ (9)
$\frac{d^{2}}{dx}g(x)$ $=\mathcal{F}^{-1}[-k^{2}(\mathcal{F}g(x))].$
@@
5 Conclusions and Limitations
In this paper, we introduced DimOL as an adaptation for major models tailored to solving high-dimensional PDEs. These include Filter-based models (FNO and CNO) and Transformer-based models such as LSM and GNOT. We discovered that dimensional awareness not only makes sense for human physicists but also machine learning models, especially when the task involves actual physical quantities as input. Our research indicates that this conclusion applies across a wide range of models and datasets. By enhancing the model’s symbolic meaning, DimOL models are better suited for reverse engineering to uncover terms with potential physical significance. While we currently use ProdLayer for DimOL, we believe that DimOL is more than this. It should be regarded as a new “dimension” to consider when designing AI4Sci models.
Notably, fully quantifying and interpreting the profits of considering dimensional awareness as a property of a model remains an open problem. One approach is to apply the model to simpler, more comprehensible datasets, such as the evolution operator for the Burgers equation in Section 4.8. This allows for identifying which dimensions are associated with the ProdLayer by analyzing the Fourier weights. However, for more complex datasets, the relationships among the Fourier weights may not have clear symbolic interpretations, making it difficult to analyze their internal connections. Therefore, it is more appropriate to view dimensional awareness as a novel and effective design principle for neural operators, rather than as a definitive scientific claim about a model’s capacity to perceive dimensions.
Another potential limitation of DimOL is its effectiveness in tasks where the input data lacks clear dimensional information or where the relationship between dimensions is highly complex or nonlinear. In these cases, DimOL may not provide substantial improvements.
Impact Statement
The potential applications and social impact of deep learning-based PDE Neural Operators (including DimOL) are vast, as they can revolutionize many fields that rely on solving complex differential equations. By leveraging deep learning to solve PDEs, these models can significantly accelerate scientific discovery, improve engineering designs, and enhance decision-making in real-world systems. Particularly for DimOL, it can universally enhance the performance of Neural Operators. By incorporating dimensional awareness into both Filter-based models (such as FNO and CNO) and Transformer-based models (including LSM and GNOT), our approach enables these models to more effectively capture the underlying physical relationships. Through the analysis of Fourier components’ weights, DimOL can symbolically discern the physical significance of each term in the model. This is exemplified in the case of the Burgers equation, where the model can identify terms corresponding to derivatives. This interpretability is crucial as it allows physicists and engineers to better understand how neural networks model physical phenomena.
While learning-based PDE solvers have the potential to accelerate and enhance simulations of complex physical systems, there are potential negative social impacts to consider. These issues largely stem from data bias, the way these methods could be deployed, and ethical concerns related to their use. For example, if the training data is skewed or incomplete (for instance, if it predominantly comes from specific environments, regions, or conditions), the model could generate inaccurate or overly generalized predictions for real-world applications.
Appendix
Appendix A Quantitative Generalization Results on New Boundary Conditions
The experiment results for Section 4.7 are recorded in Table 7.
Table 7: Prediction MSE (×$\bf\times 10^{-2}$) on the CylinderFlow dataset under different boundary conditions We have $3$ types of cylinders: Circle, Square, and Triangle. We train the models on four tasks: 1C (one circle), 1S (one square), 1T (one triangle), and ST (one square and one triangle). Therefore, CS and CT represent OOD test cases.
Model A B C BC AB AC Gain (In-Set) Gain (OOD)
FNO 4.358 5.720 4.885 12.93 12.16 11.82 - -
FNO + ProdLayer 3.998 5.403 4.622 12.38 11.85 11.16 4.488% 4.016 %
T-FNO 4.059 5.231 4.501 12.23 11.55 10.91 - -
T-FNO + ProdLayer 3.881 5.192 4.472 0.1194 11.28 10.17 2.156% 4.505%
Appendix B Hyperparameter Sensitivity Analyses
@@
Table 8: Hyperparameter sensitivity analyses on the choice of $p$ used in ProdLayer. Results are presented in MSE (×$\times 10^{-3}$). For TorusLi, we set the prediction horizon to $T=10$. For TorusVisForce, we conduct the few-shot experiments with $N=100$ and $T=4$.
Dataset Base Model Base Result $p=0$ $p=1$ $p=2$ $p=4$ $p=8$ $p=16$